OpenStudy (anonymous):

1.) Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. 2.) Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

Oh, sorry about that, Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 f(n) = 35n

OpenStudy (atlas):

There are two formulae for f(n) here. Lets try to use the first formula to find f(40) and see what happens: f(n) = f(1) + f(n-1) => f(40) =f(1) +f(39) so to know f(40), we need to know f(1) [which is given f(1) =35] and f(39) [which is not given] To find f(39), you can use the same formula again f(39) = f(1)+f(38) again we need to find f(38)!! and so on this continues.......

OpenStudy (atlas):

The second formula is simple: f(n) = 35n f(40) = 35X40

OpenStudy (anonymous):

as the differences go up by 1 each time, it is a quadratic function

OpenStudy (atlas):

If you analyse the first formula - it is actually same as the second one and gives the same result f(40) = f(1) + f(39) = f(1) + f(1) +f(38) = f(1) + f(1) + f(1) + f(37) ............................. = f(1) + f(1) +........40 times = 40 X f(1) = 40 X 35 (because f(1)=35)

OpenStudy (anonymous):

if i am not mistaken, if you start with \((0,5)\) you get \[y=\frac{1}{2}x^2+\frac{1}{2}x+5\]

OpenStudy (atlas):

The sequence is 5,6,8,11 and so on I can write the sequence as t1,t2,t3,t4 and so on where these represent the terms of the sequence t2 -t1 = 1 t3-t2 = 2 t4-t3 =3 ........... tn -tn-1=n-1 Adding up all these tn - t1 = Sum of 1st n-1 numbers = n(n-1)/2 tn = t1 + n(n-1)/2 = 5 + n(n-1)/2 So function is f(n) = 5 + n(n-1)/2